OpenStudy (anonymous):

If the hypotenuse of a 30-60-90 triangle has length 15, then the shorter leg has length

OpenStudy (anonymous):

shortcut let the lenght of shorter leg is x then x = 15sin30 = 15(1/2) =15/2

OpenStudy (anonymous):

the answer is 15/2

OpenStudy (anonymous):

or 7.5

OpenStudy (anonymous):

you can write any of these two answers